OpenStudy (anonymous):

Find the product. Simplify. 2/5 of 15 a. 5 b. 17/5 c. 6 d. 1/6

OpenStudy (anonymous):

c. 6

OpenStudy (anonymous):

how did you get this answer?

OpenStudy (anonymous):

multiply 2/5 by 15 => 15=3*5 therfore , 15/5=3 then 2*3=6

OpenStudy (anonymous):